Efficient Compilation of Forall Statement with Runtime Support

Xiaoming Li^{∗}, Yuhong Wen

Parallel Software System Group in NPAC at Syracuse University, 111 College Place, Syracuse, New York, 13244-4100

October 8, 1996

Abstract

This is an implementation document, intended as a definitive guidence for compiling FORALL statements in an HPF program, the core activity in our effort of constructing an HPF compiler on message passing platforms.

After introducing some background about the document, a model for one dimensional array and one index FORALL statements is presented. The algorithm for determining local DO loop bounds is introduced in section 3, followed by a section on communication detection and generation. Section 5 is a discussion on generalization of previous result to multi dimensional arrays and multi indices FORALL stetements. In the appendix, a study on so-called strided-subintervals is given, which serves as theoretical foundation for the techniques presented in section 4. We also provide the algorithms (as currently implemented and used by our compiler) on communication detection and coefficient calculation in appendix C and D.

∗Visiting from Harbin Institute of Technology of China